OpenStudy (yacoub1993):

Determine the solution of the system. y = x + 6 y = 2x

OpenStudy (yttrium):

You can subtract the 1st eq with the second one to solve for your x.

OpenStudy (anonymous):

If y=2x write : 2x=x+6 Got it ?:)

OpenStudy (anonymous):

@user917:Thanku :)BUTNever give the answers :)

OpenStudy (anonymous):

@Yacoub1993:Where are u ??? Please answer :(

OpenStudy (user917):

Oh, sorry @E.ali I will delete it

OpenStudy (anonymous):

No problem :) Thank u :)

OpenStudy (yacoub1993):

i am here @E.ali

OpenStudy (anonymous):

OK! Did u got that ?:)

OpenStudy (yacoub1993):

If y= 2x 2x = x + 6 2 x – x = 6 X = 6

OpenStudy (yacoub1993):

is that correct

OpenStudy (yttrium):

Yes @Yacoub1993

OpenStudy (anonymous):

@Yacoub1993 :It was right but did u got it or just got the answer ????:)

OpenStudy (yacoub1993):

i actually calculated it ad btw from where will i get the answer @E.ali

OpenStudy (anonymous):

Do u want to learn it too ?:)

OpenStudy (yacoub1993):

learn wat?

OpenStudy (anonymous):

did u understand it too ??:)

OpenStudy (yacoub1993):

yes i just need to combine both the equations

OpenStudy (anonymous):

y=2x u say that in question ... OK ?:)

OpenStudy (anonymous):

@Yacoub1993

OpenStudy (yacoub1993):

ok

OpenStudy (anonymous):

Now if we have : y=2xand y=x+6 so can write:  OK ?:)